OpenStudy (anonymous):

How do you find an equation of the line that passes through the point and has the indicated slope m? (2,4); m = -1

OpenStudy (amistre64):

y = mx -mPx + Py y = -x +2 +4 y = -x +6

OpenStudy (anonymous):

thank you so much